OpenStudy (anonymous):

What is the sum of a 7-term geometric series if the first term is 6, the last term is 24,576, and the common ratio is 4?

OpenStudy (anonymous):

-32,766 -19,662 19,662 32,766

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

Sum of a geometric series is given by: \[\huge S_n= \frac{a(1-r^n)}{1-r}\] \[\huge S_7= \frac{6[(1-4)^7]}{1-4}\] \[\huge S_7= \frac{6[-16383]}{-3}\] \[\huge S_7= 32,766\]

OpenStudy (anonymous):

thnx

OpenStudy (mathmate):

Note: \(\huge S_7=6\frac{(1-4^7)}{(1-4)}\) The final answer is correct.